# Objective-Free Local Learning and Emergent Language Structure in Thinking Machines

### v1

@@ 5.5.1. Plasticity, Emergence, and the role of Space & Time
This type of plasticity is not present in models which learn via optimizing a global objective (back propagation). While such models are typically modular, their architecture is fixed beforehand. Though the black box paradigm permits a type of data-driven “emergence”, it is not the type of structural emergence described here, but instead reflects a discovery of internal representations allowed within the fixed architecture. The attention mechanism of an LLM is assumed beforehand (Vaswani et al., 2017), not a learned structured. This differs from Conway’s Game of Life, which can “learn” a Turing machine – not from recognizing patterns in data nor minimizing an objective, but as an accident due to the local dynamics & initial conditions (Rendell, 2011, 2014).
The hierarchical key-value memories described here can be understood as arising due to the interplay of space & time. The spatial aspect is reflected in how the key-value relationship preserves the morphological structure, such that the value part encodes how to complete a word from a triggered key. Thus no positional embeddings are necessary (or helpful). This spatial awareness is learned because retokenization makes replay boundary aware; so that the added embedding neuron correlates with the features corresponding to smoothly correlated regions of string. By having the neuron activated throughout this cycle, it learns this smooth feature set as if all its elements fired simultaneously during a single Hebbian update step. At face value this process is “precisely timed”, but thanks to the spatial awareness of retokenization, doing this precise timing is trivial.
It’s useful to think of replay as a minimally structured game (or a local dynamic coding scheme) – grab a new neuron, poke short-term memory to randomly replay its correlations, then repeat the process again with a new neuron once replay terminates. It works because the local event learning makes continued learning & emergence incidental. The assumed structure of replay is chosen to exploit this “accidental learning”. Likely other (more efficient less random) structured games exists which exploit this accident.
Even following our minimal replay blueprint, the number of possible different emergent structures grows combinatorially with the number of learned long-term embeddings. As more embedding neurons populate the long-term memory ($V_{\Omega}:=(v_{k},a_{\alpha},b_{\xi},\cdots)_{\Omega}$), the model can learn correlations between these embeddings ($V_{\Omega}V_{\Omega^{\prime}}$ for subset of $\Omega$), which can then form entirely novel compositional structures (beyond hierarchical & super-hierarchical). These newly learned structures, in turn, enrich the token representations available to short-term memory, enabling further complexity. Classifying all possible emergent structures, given the local dynamics, is akin to exactly solving a physics problem. In general this is not always possible, but nevertheless opens a potentially fruitful direction for future research.
Figure 3. Hierarchies of hierarchies: STM’s are n-gram models which build features ($\tilde{v}$,$\tilde{a}$,$\tilde{b}$,…) from relevant correlations ($\tilde{v}v$,$\tilde{a}a$,…). These can be replayed into symbolic key-value memory tokens of increasing complexity LTM ($a$,$b$,…). We show the emergent inference classes learned at each step. In addition to the direct hierarchies, LTM tokens of mixed class can be correlated.
6. Toy Example: Finger Counting
This example illustrates how replay and hierarchical memory enable the model to bind structured inputs (e.g. finger gestures) to symbolic tokens (e.g. number words).
6.1. Modalities and Embeddings
We assume two input modalities:
  • Letter modality: Basis tokens $v_{k}$ for alphabet letters (e.g., $v_{\bf o},v_{\bf n},v_{\bf e}$).
  • Finger modality: Scalars $u\in\{0,1\}$ indicating raised fingers.
Each embedding $a_{\alpha}$ is associated with hierarchical memory:
$m_{\alpha}=\sum_{n}\sum_{\mu_{n}}m^{(n)}_{\alpha,\mu_{n}}\tilde{v}^{(n)}_{\mu_{n}}$
For example, the word “one” is stored as:
$m_{1}$ $=$ $m_{1,\bf o}^{(1)}v_{\bf o}+m_{1,\bf n}^{(1)}v_{\bf n}+m_{1,\bf e}^{(1)}v_{\bf e}$
$+$ $m_{1,\bf on}^{(2)}\tilde{v}_{\bf on}^{(2)}+m_{1,\bf ne}^{(2)}\tilde{v}_{\bf ne}^{(2)}+m_{1,\bf one}^{(3)}\tilde{v}_{\bf one}^{(3)},$
and similarly for “two” and “three”.
Finger gestures are retokenized as:
$\tilde{u}^{(n)}=\prod_{x=1}^{n}u_{x},$
with $\tilde{u}^{(n)}=1$ if $n$ fingers are raised.
6.2. Binding Fingers to Numbers
We define the Hamiltonian:
(41) $H$ $=\sum_{\alpha}a_{\alpha}\left(\sum_{n}\sum_{\mu_{n}}m^{(n)}_{\alpha,\mu_{n}}\,\tilde{v}^{(n)}_{\mu_{n}}\right)$
(42) $\quad+G_{1}^{(1)}\,a_{1}\,u+G_{2}^{(2)}\,a_{2}\,\tilde{u}^{(2)}+G_{3}^{(3)}\,a_{3}\,\tilde{u}^{(3)}.$
where the $G$ couplings are learned via Hebbian updates, e.g. showing two and two raised fingers. (Technically, replay will bind $\alpha>n$ embeddings to $\tilde{u}^{(n)}$, such as $a_{3}$ being bound to $u$, $\tilde{u}^{(2)}$, and $\tilde{u}^{(3)}$. But this interference is mitigated by biasing $G_{\alpha}^{(n)}\rightarrow\beta^{n}G_{\alpha}^{(n)}$ with $\beta>1$, and choosing $f_{\alpha}$ to be normalization.)
During inference we raise 2 fingers ($\tilde{u}^{(n)}=\sum_{n^{\prime}\leq 2}\delta_{n^{\prime},n}$),
$\Delta a_{\alpha}=G_{\alpha}^{(n)}\tilde{u}^{(n)}\quad\Rightarrow\quad a_{\alpha}\mapsto\text{``two''}\Rightarrow v_{k}=t,w,o.$
That is, a finger count activates the corresponding embedding, which then drives autoregressive symbol generation following the key-value relationship encoded in Eqn 41.
7. hierarchical feature mixing & random language
We begin by analyzing the structure of a randomly sampled vocabulary, consisting of 100 personal names (in English or Spanish) (dat, aset). For this set, we plot the $d_{n}$ distribution, where $d_{n}$ counts the number of unique $n$-grams observed at each scale. The resulting $d_{n}$ distribution follows an approximately log-normal form, with the largest $n$ determined by the longest name in the sample.
In contrast, cutting closer to $n_{\text{peak}}$ loosens the scale-dependent constraints, allowing novel $n$-grams to form. Nevertheless, even here, the surviving short-scale DAG structure continues to guide regrowth, ensuring that generated words remain smooth and morphologically coherent.
We find that though the original notion of smoothness is perfectly preserved below the cut, cutting away more of the $d_{n}$ peak causes the hierarchy to begin losing its recognizable English/Spanish nature. This highlights the importance of the $n_{\text{peak}}$ grams to language structure [^5] For $n\lesssim n_{\text{peak}}$, $d_{n}$ grows approximately exponentially, with learned constraints ($g^{(n)}=0$) pruning branches and thereby reducing the effective growth rate. For $n>n_{\text{peak}}$, the smoothness condition takes over: new tokens are permitted only if all subgrams are valid, which recursively restricts composition and causes a collapse in $d_{n}$. This collapses guarantees finite word length (Eugenio, 2025). In formal language terms, this might be described as a consequence of the regular-language-like limitations of stored strings..
Finally, if we cut at $n_{\text{cut}}=2$, we discard all memory of the original names. This is the data-free regime, where growing a hierarchy generates what we refer to as a “random language” (Eugenio, 2025). Nevertheless, growing hierarchies under the smoothness constraint still yields finite word-lengths and an approximately log-normal $d_{n}$ distribution (HERDAN, 1958; WILLIAMS, 1940; Limpert et al., 2001). This shows that bounded, hierarchical morphology is an intrinsic feature of the model itself, independent of any training data, and that language structure arises naturally from smooth hierarchical learning.
7.1. Synaptic Zipf-laws
Random languages generated by our model exhibit rank-ordered $n$-gram frequency distributions that follow a scale-invariant Zipf-like law (Zipf, 1949; Gan et al., 2009), a signature of morphological structure (Eugenio, 2025). These spectra provide insight into the statistical relevance of subword patterns, ranked from most to least frequent, and reveal how this structure scales with increasing pattern length.
Note we can interpret these ranked frequencies as defining an effective memory spectra $g^{(n)}_{\text{eff}}$, where Hebbian learning imprints frequency counts from local events $\mathcal{E}_{n}$ directly into memory:
$g^{(n)}_{\text{eff}}=\sum_{\mathcal{E}_{n}^{\small\text{vocab}}}\tilde{v}v.$
In the absence of forgetting (Eugenio, 2025), this spectrum reflects all observed correlations of length $n$.
Fig 4 shows the effective model spectrum corresponding to a model whose memory graph was grown randomly under the hierarchical DAG constraint. As such, the resulting distribution reflects only the inductive bias of the memory structure itself.
At short lengths ($n<n_{\text{peak}}$), constraints are minimal, and the distribution appears structureless: a Zipf-like power-law head followed by a flat tail of irrelevant $n$-grams, similar to those found in uniform random strings (Eugenio, 2025; Gan et al., 2009). However, as $n$ increases, the DAG constraint increasingly shapes memory formation. Beyond the peak, the tail vanishes and the power-law scaling persists, signaling the emergence of a dominant compositional structure at longer lengths.
Figure 4. The frequency an $n$-gram (length $n$) appears in the vocabulary, rank-ordered from most to least frequent. The vocabulary is for the random language bjfyi (in supplement). Equivalently, this plot can be understood as representing an effective model spectra ($g^{(n)}_{\text{eff}}$ per $n$). Scale-invariance emerges beyond $n_{\text{peak}}=3$ due to the compositional structure.

### v2

@@ 6.5.1. Plasticity & Emergence
This type of plasticity is not present in models which learn via optimizing a global objective (back propagation). While such models are typically modular, their architecture is fixed beforehand. Though the black box paradigm permits a type of data-driven “emergence”, it is not the type of structural emergence described here, but instead reflects a discovery of internal representations allowed within the fixed architecture. The attention mechanism of an LLM is assumed beforehand (Vaswani2018_AttentionAllYouNeed), not a learned structured. This differs from Conway’s Game of Life, which can “learn” a Turing machine – not from recognizing patterns in data nor minimizing an objective, but due to the local dynamics & initial conditions (PaulRendell2011_ConwayTuringComplete; PaulRendell2014thesis_ConwayTuringComplete).
It’s useful to think of replay as a minimally structured game (or a local dynamic coding scheme) – grab a new neuron, poke short-term memory to randomly replay its correlations, then repeat the process again with a new neuron once replay terminates (Fig 7(a)). While structured, what information is replayed is entirely random, and thus no non-local dependence of specific information stored in those memories is required. This should not be confused with “replay” as defined elsewhere in the machine learning literature, which involves replaying specific memories in order to prevent their forgetting in a fixed architecture (Parisi2019review_Forgetting&Replay). Here replay generates new feature basis in a non-static architecture—nothing is ever written over, as the memory capacity is unbounded, and thus there is no catastrophic forgetting.
Figure 8. Hierarchies of hierarchies: Short-term memory structures are retokenizing n-gram models which grow feature tokens ($\tilde{v}$$\tilde{a}$$\tilde{b}$,…) from relevant correlations (products of tokens: $\tilde{v}v$$\tilde{a}a$,…). These can be replayed into symbolic object tokens of increasing complexity ($a$,bb,…). The simplest object tokens (first-order embeddings) decode as a key-value memory (temporal association via simultaneity) wrapping an n-gram model (captures spatial structure). Not shown: mixed correlations (e.g $ab+ba$) can also be replayed into tokens.
7. Simplifications & Code
7.1. One-memory-per-retokenizer limit
Up to this point, we have treated the problem as one of: grow short-term memory, then disentangle it into individual long-term memory tokens. However, if the boundaries in the data are known, then its possible to skip the disentanglement step.
In this regime, each retokenizer is trained on a single input pattern, such as one word or sequence. The local Hebbian updates (Eqn 7) therefore construct a minimal hierarchical graph whose projectors $\{P_{n}\}$ collectively describe a single smooth path terminating in one node of the DAG. All sub-grams of the input are locally merged according to the smoothness constraint, yielding an internally consistent projector chain $P_{N}P_{N-1}\cdots P_{2}$ representing that single stored memory.
The corresponding embedding is then formed directly by association (i.e direct sum) following Eqn 27. For the rest of the paper, we will simplify our notation, and simply write the embeddings like
(47) $a_{\bf cat}=(v_{\bf c}+v_{\bf a}+v_{\bf t})\oplus(\tilde{v}^{(2)}_{\bf ca}+\tilde{v}_{\bf at}^{(2)})\oplus\tilde{v}_{\bf cat}^{(3)},$
[…]
7.2. Boundaries are data
Note that higher-order embeddings (Sec 6.5) can be constructed similarly. For example,
(48) $b_{\bf the\ cat\ jumps}=$
$(a_{\bf the}+a_{\bf cat}+a_{\bf jumps})\oplus(\tilde{a}^{(2)}_{\bf the\ cat}+\tilde{a}_{\bf cat\ jumps}^{(2)})\oplus\tilde{a}_{\bf the\ cat\ jumps}^{(3)}$
Again, this requires specifying the statistical boundaries, such as specifically preparing small phrases or sentences. Thus there is an arbitrariness in learning from structured data, namely the structure itself: boundaries are data. This includes boundaries not necessarily indicated by notation in writing. If not made explicit, the model will construct its own interpretation of the boundaries, which may not align with the intended behavior, but will align with human language universals.
Note that the coordination length grows exponentially with a linear increase in token depth (how many times high-order tokens were correlated & replayed to build yet even higher-order tokens).
7.3. Code & Latent Space Topology
Self-organizing neural learning is inherently multi-modal: learning new embeddings expands it computational substrate. So far we have used $v$ for external or input modalities and ($a,b,\ldots$) for internally generated ones. By default, all tokens are treated as being part of a giant undifferentiated content-addressable library. Thus for simplicity, we will unify internal tokens under the $a$-notation.
The accompanying code (ababot2025) follows this structure directly. It is organized around high-level Token objects, each providing built-in methods Token.key() and Token.retrieve(). Collections of tokens are grouped in Pool objects, which share the same interface but represent multi-token bases used by the retokenizer to learn correlations. A Pool thus acts as a composite token defining the active basis for local learning.
Consider the embedding from the previous section: $a_{\bf the\ cat\ jumps}$. It has three children: $a_{\bf the}$, $a_{\bf cat}$, & $a_{\bf jumps}$. These children define the basis Pool used to construct it. Decoding the latent token requires keeping track of both the projectors
(49) $P_{2}$ $=$ $\tilde{a}^{(2)}_{\bf the\ cat}a_{\bf the}a_{\bf cat}+\tilde{a}^{(2)}_{\bf cat\ jumps}a_{\bf cat}a_{\bf jumps}$
(50) $P_{3}$ $=$ $\tilde{a}^{(3)}_{\bf the\ cat\ jumps}\tilde{a}^{(2)}_{\bf the\ cat}a_{\bf jumps}$
and the three child tokens themselves.
The code handles token tracking via two complementary storage modes: inline and registry builds. The latter exploits a global token registry to keep track of the parent-child relationships. This provides a lightweight global store, allowing a Token to carry only a reference (TokenRef) and resolve its parameters dynamically at retrieval.
Alternatively, the inline mode stores a token and its children directly together (as opposed to storing a pointer to the children, as with the registry). If the children themselves have children, such as with our example, then all descendants are stored together under the major token.
The inline mode has the effect that no child token is shared by multiple parents, as each carries its own degenerate copy of its children. While this requires more disk space, only the part of the inline token necessary for calculation needs be stored into working memory. It is therefore not less memory efficient, and has the benefits of each token being independent from all other tokens.
7.3.1. Token modes (GI vs. Pnn).
The code supports two Token modes for representing the projector structure: a (default) gauge-invariant GI mode, and an ultra-compressed Pnn mode.
The default mode treats a single set of projectors $\{P_{n}\}$—one per DAG depth $n$. It allows for exact retrieval & left/right inference.
The Pnn mode follows the SVD-based matrix product compression of Sec 6.4. In this mode, each DAG depth gets its own depth-specific projectors, such that Token memory is organized as
$\texttt{Token.mv\_alpha}=\begin{bmatrix}[m^{(1)}]\\ [P_{2,2}\!,\;m^{(2)}]\\ [P_{3,2}\!,\;P_{3,3}\!,\;m^{(3)}]\\[-3.0pt]$ $\vdots\end{bmatrix}$
which do not form a gauge-invariant set. The reasons for this are discussed in Sec 6.4. Note however, the compression is still exact, such that the complete pattern can still be retrieved (even without perfect tokenization). This makes it technically possible to convert between Pnn & GI via replaying Pnn back into an empty retokenizer.
In the current implementation, we instead choose to exploit a practical gauge hack: summing left/right gauges, $P\!\leftarrow\!P\oplus P_{\mathrm{Left}}$. This merges right & left inference paths to a single terminal node, and converts right-token prediction into an “outward-token” prediction, which generates either a left or right token. Entropy is used to estimate boundaries & determine which side to merge the token.
8. Trees & Syntax
In summary, human language patterns at all structural levels reflect a universal mechanism of subword formation, which arises in the limit of a sub-symbolic system constrained by locality.
8.1. The Embedding Graphs are Syntax Trees
Local learning naturally produces a tree topology of embeddings (Fig 9). Each retokenizer builds correlations from leaf to root, so the resulting parent–child graph of feature tokens forms a locally organized tree. The ordering of the leaves arises directly from learned correlations, not from any predefined syntax rule. This structure reflects how global order emerges sub-symbolically in a local system.
Because the embedding graph is tree-structured and explicit, its internal organization is interpretable. The same topology that emerges from locality mirrors the syntax trees used in symbolic theories of language (Keenan&Moss2016book_MathematicalStructuresInLanguage). This allows symbolic structure to be passed directly to the model: large classes of word embeddings can be tied by simple association to form composite trees.
Such “soft-coded” initialization is effectively a type of pre-training, where we exploit the natural hierarchical organization to directly grow a symbolic theory as a topological enforced memory.
8.2. Subregular Symmetry
Note that hierarchy isn’t handled via symbolic rules, but is an implicit feature of how the memory is organized. Hierarchical boundaries are statistical boundaries, and hence sub-symbolic. This is the same reason why there are no space tokens (see Sec 3.3.2).
Because the retokenizer is local, it cannot distinguish a latent representation from a letter token. This makes it so that all ordered patterns, at any level, exhibit universal sub-word structure. This subregular symmetry mirrors statistical observations in human language (Herdan1958_logNormal_intraword; Williams1940_logNormal_interwords) and recent formal work unifying syntax under subregular grammars (Graf2022_SubregularLinguistics; GrafHanson2025_SubregularSyntax).
8.2.1. Memorization: feature vs object tokens
Note that there is an arbitrariness in representation arising from where one chooses to place the statistical boundaries (e.g Fig 9(a)). The pattern abc could be learned as an ordered 3-point retokenized pattern (i.e intra hierarchy); or inter-hierarchically, ordered as a pair of retokenized 2-point memories.
In the Figure example shown, the minimal total number of neurons required for either choice are equal. However, for patterns larger than 3, it becomes more memory efficient to learn inter-hierarchically. This is because retokenized patterns require learning all sub-patterns as feature tokens. [^16] This is also true on the computer: larger patterns means wider projectors. But any pattern can be stored inter-hierarchically, so that technically no tensor need have a dimension size larger than 2. This amounts to being able to trade width for depth (more but smaller tensors).
This aligns with understood limitations of short-term memory (Shiffrin1970_forgettingLongStrings) (the retokenizer being a compositional short-term memory structure). If the reader is asked to memorize alsjfgauy, they might have an easier time by first breaking it up into als, jfg, & auy. This being a common tactic for remembering long numbers.
8.3. Emergent Categories
The same mechanism that builds meaningless subword structure also enables the self-organization of categories through replay. In key-match replay, a random latent token is decoded from long-term into short-term memory. If the token corresponds to a word such as cat, all embeddings that most closely overlap with it (e.g. the cat, a cat runs, cat jumped) are replayed together. Their entangled feature maps can then be bound to a new embedding, forming a bookmark for future key searches and reducing the search space to similar contexts. These bookmarks act as new internal modalities—feature bases representing contextual associations—that can themselves be correlated and replayed to generate higher-order symbolic structure. We explore these emergent replay behaviors and their role in category formation in a follow-up work.
(a) Two examples demonstrating how to formally map between embedding & tree structure. There is an arbitrariness in representation arising from where one chooses to encode the boundary information.
(b) Categories can be formed by associating relevant tokens. Above shows a simple (determinant-noun)-verb construction.
Figure 9. Local Tree Topology of Long-Term Memory
9. Toys
We present a couple toy models to highlight model interpretability, as well as qualities special to it being a neuro-symbolic bridge.
9.1. Latent Codes
We start with the lower-case English alphabet as our basis, and grow embeddings $a_{\alpha}$ corresponding to an arbitrary list of words (cat,dog,fish,gator,…). We then associate all words into new class neurons (labeled A,B,C), organized by word length,
(51) $a_{\bf A}$ $=$ $a_{\bf cat}\oplus a_{\bf dog}\oplus\cdots$
$a_{\bf B}$ $=$ $a_{\bf fish}\oplus a_{\bf cake}\oplus\cdots$
$a_{\bf C}$ $=$ $a_{\bf gator}\oplus a_{\bf smith}\oplus\cdots$
Note that patterns in the latent space, e.g $a_{\bf A},a_{\bf C},a_{\bf B}]$, behave like codes, in that the subsequent decoding of tokens in the list produces one possible valid string satisfying that order.
Step-by-step, we activate $a_{\bf A}$, which activates all length-3 words. We then sample from that subset, generating (say) $a_{\bf cat}$, which decodes via the DAG structure of Eqn 47 into precisely $v_{\bf c},v_{\bf a},v_{\bf t}]$. Fully decoded gives us something like, “cat smith cake”—one of many possible 3-5-4 ordered strings.
We can store the latent code as the token
(52) $a_{\bf ACB}=(a_{\bf A}+a_{\bf B}+a_{\bf C})\oplus(\tilde{a}_{\bf AC}^{(2)}+\tilde{a}_{\bf CB}^{(2)})\oplus\tilde{a}_{\bf ACB}^{(3)}.$
Even more complex patterns can be orchestrated as higher-order latent codes.
Note that we could instead have formed our vocabulary randomly, and simply formed random associations of the random vocabulary. A tree structure corresponding to multi-scale symbolic order naturally forms from these random processes. Such random structures are not content-less, as embeddings have meaning arising from their relative information. It has content, it’s just random content.
9.2. Quantity Tokens
Consider a 1-letter alphabet ($d=1$). We’ll write the basis token for this mode as $u(x)$, which is $1$ (or 0) if the $x^{\text{th}}$ element in the string is the letter (or empty space). Its corresponding feature tokens $\tilde{u}^{(n)}$ measure quantity. They are formally defined from scalar generalizations of the projectors, which take the value $1$ or 0 depending on if the trailing $n$ tokens aren’t empty. Likewise, quantity object tokens can be defined
(53) $q_{n}=\frac{u\oplus\beta\tilde{u}^{(2)}\oplus\cdots\oplus\beta^{n-1}\tilde{u}^{(n)}}{\beta^{n-1}},$
where $\beta>1$ biases toward longer length scales. This makes $q_{n}$ behave like $\tilde{u}^{(n)}$, but with a small residual overlap (vanishing in $\beta\rightarrow\infty$ limit) with all $\tilde{u}^{(n^{\prime})}$ with $n^{\prime}<n$.
Quantity tokens are a consequence of grounding symbolic behavior in sub-symbolic estimation. When $q_{n}$ & $q_{m}$ are decoded in sequence, the resulting combined [^17] Decoding is boundary aware, such that we can decode $q_{1},q_{2}]$ as $[u,0,u,u]$ or $[u,u,u]$, depending our choice to keep/throw away boundary information. sequence keys $q_{n+m}$. Thus they provide the neural net an internal metric. Embeddings associated with $q_{n}$ inherent its quantity $n$.
9.3. Feature Shadow & Inter-Modality
In addition to the $u$ tokens, we’ll introduce the lower-case English alphabet ($d=26$), and associate every letter with $u$:
(54) $a_{\bf a}$ $=$ $v_{\bf a}\oplus u$
$a_{\bf b}$ $=$ $v_{\bf b}\oplus u$
$\vdots$
$a_{\bf z}$ $=$ $v_{\bf z}\oplus u$
The new internal modalities are keyed by length $1$ context in the $u$ channel. Rather than treat them as separate channels, we can define the map
(55) $u(x)=\theta\Big(\sum_{j}v_{j}(x)-1\Big),$
which is a step-function, acting on every element of the $v$ string at position $x$ in the context. Thus $u(x)$ lights up if any of the $26$ neurons of $v_{j}$ at site $x$ fire.
Now any higher-order token learned from these modes inherent a notion of length equal to the length of their stored pattern. For example,
(56) $a_{\bf aba}=(a_{\bf a}+a_{\bf b})\oplus(\tilde{a}^{(2)}_{\bf ab}+\tilde{a}^{(2)}_{\bf ba})\oplus\tilde{a}^{(3)}_{\bf aba}$
is keyed by context $[u,u,u]$. As are all other length $3$ word embeddings in memory.
Now consider the following list of latent tokens: $q_{3},q_{5},q_{4}]$. Fully decoded into $u$, it looks like
$[u,u,u,0,u,u,u,u,u,0,u,u,u,u].$
The regions between the 0’s project up into $\tilde{u}^{(n)}$, and therefore overlap with all tokens storing patterns of that length. This occurs even though $q_{n}$’s are defined independent of $a_{\alpha}$’s, and is a consequence of their shared projected behavior, not any direct correlation or shared association [^18] Note that the feature tokens $\tilde{u}^{(n)}$ of both $q_{n}$ and $a_{\bf aba}$ aren’t necessarily literally the same. Feature tokens $\{\tilde{u}^{(n)},\tilde{v}^{(n)},\tilde{a}^{(n)}\}$ are better understood as being transient placeholders representing how context is projected. The projectors don’t have to be identical (due to representational freedom of dense matrices), so long as they behave equivalently. In the code, each token stores its own independent projectors..
This pattern of quantities can be stored as the token:
(57) $a_{3,5,4}=(q_{3}+q_{4}+q_{5})\oplus(\tilde{q}_{3,5}^{(2)}+\tilde{q}_{5,4}^{(2)})\oplus\tilde{q}_{3,5,4}^{(3)}.$
Note that because we have multiple external modes ($v_{k}$ & $u$), there is a freedom in how to decode. Step-by-step, activating $a_{3,5,4}$ drives (say) $q_{4}$ to fire, which decodes out to its boundaries as $[0,u,u,u,u,0]$. If before decoding further in $u$-channel, we stop and form a list (a.k.a a Pool or temporary association) of all embeddings which overlap with that context, then randomly sampling from that list behaves like sampling from a list of length-4 words. More generally, by choosing a decoding order for the input modes ($u$ then $v_{k}$), we can decode governed by the inter-modal relationships. Then fully decoded $q_{3,5,4}$ looks like $a_{\bf ACB}$ from the previous subsection.